# Sigma: Scalable Spectral Insights for LLM Collapse

### v1

@@ 1 Introduction
LLM Collapse (shumailov2024; Shumailov2024-dj) refers to the phenomenon where Large Language Models (LLMs) progressively degrade towards a degenerate state. This state arises as a result of recursive training, where models are trained on iterations of data generated by other models, often previous versions of themselves. We term this model-generated output Synthetic Data, in contrast to organic data produced by human users.
@@
In the current landscape of AI research, the demand for reliable and high-quality training data has never been higher. The scale of the latest generation of LLMs is increasing at such a staggering speed that the natural generation of internet content by humans cannot keep pace. Consequently, utilizing synthetic data has become a vital, if not inevitable, step in LLM training. While this is a logical progression, it carries a significant caveat. Without the infusion of new data produced by human users, model collapse occurs as the model iterates on its own probability distribution. Ultimately, the model may converge to a state characterized by low variance and degraded representational quality (shumailov2024; Shumailov2024-dj).
@@
Therefore, it is imperative to conduct a rigorous mathematical examination of LLM collapse. What are the underlying mechanisms at the mathematical level? How can this degradation be tracked and quantified throughout the training process? Furthermore, how can we leverage these metrics to gain deeper insights into LLM dynamics and guide the development of more efficient training architectures? This paper aims to answer these questions.
@@
We introduce Sigma (Spectral Inequalities for Gram Matrix Analysis), a novel framework that quantifies and benchmarks model collapse by utilizing the spectral bounds of the Gram matrix of LLM embedding vectors. The Gram matrix is a classical mathematical concept named after Jørgen Pedersen Gram. It has a wide range of applications in fields such as Riemannian geometry and control theory, and has found critical utility in modern applied mathematics and computer science domains, including the finite element method (10.5555/894170). It is also widely used in classical NLP as an intuitive method to encode variable quantities of natural language embeddings into a fixed-dimensional matrix representation(10.1162/153244302760200687; NIPS2014_b7866697). As we enter the era of LLMs, the Gram matrix remains a vital tool, as it captures the intrinsic geometry and distributional properties of the data we seek to analyze. Sigma revisits this fundamental operator to quantify the modern phenomenon of representation collapse.
@@
Inevitably, the spectrum of the Gram matrix becomes the key to understanding LLM collapse. We develop our theory and framework around this central mathematical object, allowing us to monitor, estimate, and predict the behavior of LLMs regarding model collapse through training iterations. Crucially, our method utilizes both deterministic and stochastic bounds to function effectively even with partially observed data, ensuring our framework remains scalable for large-scale models.
@@ 2.1 Related Work
The risk of training generative models on their own outputs has been formalized as model collapse, where repeated self-training progressively erodes distributional support, particularly in the tails, and can culminate in highly generic or degenerate generations (shumailov2024; Shumailov2024-dj).
Related phenomena have been studied under different labels and experimental protocols, including Model Autophagy Disorder (MAD), which emphasizes precision–recall tradeoffs and the role of “fresh” real data injections across generations (alemohammad2023selfconsuminggenerativemodelsmad). Recent theory sharpens these concerns by distinguishing regimes where degradation is transient versus strong and asymptotic, even when synthetic data is mixed with a constant fraction of real data (dohmatob2024strongmodelcollapse).
Complementary analyses interpret recursion as an entropy- or variance-contracting process and propose tail-centric quantities such as time-to-forget and covariance shrinkage to explain why rare events vanish first (seddik2024badtrainingsyntheticdata). Other lines formalize stability of self-consuming loops via recursion-stability notions, connecting collapse to unbounded generalization degradation across generations (yoon2025modelcollapseselfconsumingchain).
Finally, practical collapse signals in text include over-concentration of surface features (e.g., $n$-grams), emphasizing the need for diagnostics that are sensitive to diversity loss beyond average quality metrics (zhu2025synthesizetextdatamodel). Most prior work studies recursion within a single-model loop (or a fixed teacher–student chain), whereas real-world synthetic data arises from multi-model, retrieval-mediated ecosystems. Building on this black-box perspective, recent work develops statistically grounded uncertainty and calibration techniques tailored to LLM-as-a-Judge systems (pang2025unsupervisedconformal; pang2025tamingvariability), and studies related dynamics that can shift judge behavior over time (wang2025llmwebdynamicstracing), where LLM Web Dynamics (LWD) explicitly targeting this setting by modeling a network of LLMs coupled through a retrieval-augmented substrate, and analyzes convergence patterns under such web-like feedback.
@@ 3.3 Error Estimation
Naturally, we will be interested in how much error does the theorems in the pervious section introduce. In fact, 2 provides extra insights for the tightness of inequalities in 1. Let $C$ be the covariance matirx defined in theorem 2. We know that for large $n_{A},n_{k}$, $\det(G^{(k)})\approx\left(n_{k}/n_{A}\right)^{m}\det(G_{A}^{(k)})\approx n^{m}\det(C)$, we can define the asymptotic Overestimation Ratio $\mathcal{R}$ as the ratio of the Weyl bound to the stochastic expected value:
@@ 3.4 Sigma-UB Collapse Monitoring
Section 3.2 establishes the theoretical foundation for detection, but the raw bounds rely on unobserved quantities ($G_{B}^{(k)}$) or unknown constants ($K$). This subsection translates those theorems into two scalar, fully computable diagnostics, computed only from $\left(G_{A}^{(k)},n_{A},n_{k}\right)$.
Regularization Convention ($\delta$). We fix a small constant $\delta>0$ and evaluate log-volumes using $\log\det(\cdot+\delta I_{m})$. The same $\delta$ is reused across all checkpoints to ensure comparability and numerical stability under rank deficiency. This regularization ensures the metric is well-defined and Lipschitz continuous with respect to small perturbations in the embeddings.
Now we turn both theorem 1 and theorem 2 into computationally scalable monitoring metrics via the following two bridges, respectively.
Diagnostic I (Theorem 1): Tail-Energy Budget. Note that we never construct $B^{(k)}$ and thus cannot measure $\beta_{k}$. Instead, we enforce a standard embedding preprocessing (e.g., $\ell_{2}$-normalization or clipping) so that $\|v_{j}^{(k)}\|_{2}^{2}\leq\rho$ for all generated responses. With $n_{2}:=n_{k}-n_{A}$ unseen columns, this yields the computable worst-case budget
@@
Appendix B shows that substituting $\widehat{\beta}_{k}$ produces a valid one-sided envelope for the full regularized log-determinant.
Diagnostic II (Theorem 2): Constant-Free Size Correction. Recall that Theorem 2 implies the full model entropy is effectively the sub-sample entropy shifted by the geometric scaling factor $m\log(n_{k}/n_{A})$. While the exact value depends on an unknown constant $K$, we can eliminate this constant by monitoring the trend rather than the absolute value.
Combining this with Theorem 1, we operationally track collapse via two complementary diagnostics: (i) a size-corrected statistic that relies on this stochastic scaling, and (ii) a normalized deterministic envelope that uses a computable tail budget.
Two Computable Diagnostics. We monitor collapse using (i) a normalized deterministic envelope and (ii) a covariance-normalized LLN proxy:
@@
$\quad-m\log(\widehat{\beta}_{k}+\delta),$ (3.16)
@@
We report baseline drifts
@@
Sigma-UB Monitoring Algorithm. Algorithm 1 computes both diagnostics at each checkpoint using only the sub-Gram matrix and counts. In practice, $\log\det(\cdot)$ is computed via Cholesky on the shifted PSD matrix (Appendix B).
@@
How to Interpret the Outputs. Both diagnostics decrease when eigenvalues of $G_{A}^{(k)}$ contract, with heightened sensitivity near the $\delta$ floor. Operationally, a sustained negative drift in either $\Delta\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ or $\Delta\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)$ signals collapse onset; saturation indicates late-stage rank collapse. The two-track design is intentional: $\Delta\mathcal{G}_{\mathrm{KF}}$ is a deterministic, conservative envelope controlled by a known tail-energy budget, while $\Delta\mathcal{U}_{\mathrm{LLN,cov}}$ is a sensitive size-corrected proxy whose unknown constants cancel under differencing. Disagreement between the tracks is itself informative, often indicating violations of the consistent-context assumption (prompt drift, retrieval drift, filtering changes) rather than numerical artifacts.
@@ 4 Experiments
4.1 Experiment Objectives and Controlled Assets
Our experiments measure the operational behavior of training with recursively regenerated synthetic text under tightly controlled, evaluation-stable conditions. We isolate two questions, aligned with two settings (S1–S2):
(1) Data recursion (S1). When synthetic text is regenerated every generation but each training run restarts from the same base checkpoint, do the SIGMA geometry diagnostics drift progressively, indicating representational contraction driven purely by data feedback?
(2) Weight recursion (S2). Under the same regeneration protocol and real:synth mixture rule, does carrying model weights across generations materially amplify or alter collapse dynamics relative to restart-from-base training?
Assets and evaluation stability. We reuse the license-clean real corpora and frozen prompt banks from DecayBench as controlled assets. Each domain
$d\in\{\textsc{CIVICS},\textsc{FINANCE},\textsc{SCIENCE},\textsc{SOFTWARE}\}$
provides: (i) a real corpus $R^{(d)}$ pre-sliced into non-overlapping 64-token contexts, and (ii) a fixed prompt bank $P^{(d)}$ with stable prompt IDs (and bucket labels). To avoid evaluation drift, all checkpoints are evaluated on the same frozen prompt bank $P^{(d)}$ with fixed decoding and a frozen sentence encoder. Following the main-text budget, we report results on a single target domain (FIN) and provide full numeric tables and auxiliary analyses in Appendix C.
4.2 Models and Fine-Tuning Setup
Unless otherwise noted, we start from a single base checkpoint $M^{(0)}$ (e.g., OPT-125M) and fine-tune with a fixed optimizer, training budget, and data formatting across all conditions. All training examples are stored as 64-token text blocks so that token budgets remain comparable. Full hyperparameters and run metadata (optimizer/LR, steps, batch size, weight decay, decoding, seeds) are consolidated in an experiment card (Appendix C.2).
4.3 Synthetic Generation, Mixture Rule, and Two Recursion Settings
@@
Regeneration protocol and constant-budget mixtures. For a fixed domain $d$, we use each real 64-token context $x\in R^{(d)}$ as a prompt and sample a fixed-length continuation from the previous-generation model to form a synthetic pool $S^{(g,d)}$ at generation $g\geq 1$. Training data at generation $g$ is a constant-budget mixture of synthetic and real examples: a fraction $(1-\alpha)$ are drawn from $S^{(g,d)}$ and a fraction $\alpha$ from $R^{(d)}$, with the total number of training blocks held fixed to $|R^{(d)}|$. The real fraction $\alpha\in[0,1]$ is held fixed within each run.
@@
Setting S1: restart-from-base (data recursion only). Each generation regenerates a fresh synthetic pool, but training always restarts from the same base checkpoint:
$M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(0)},\,D^{(g,d)}(\alpha)\big).$
This isolates collapse induced by the evolving synthetic data distribution without compounding weight updates across generations.
@@
Setting S2: true recursion (data + weight recursion). This setting carries weights across generations under the same regeneration and mixture rule:
$M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(g-1,d)},\,D^{(g,d)}(\alpha)\big).$
Comparing S1 vs. S2 therefore isolates the incremental effect of weight recursion under a matched synthetic regeneration pipeline.
4.4 Collapse Measurements and Baselines
Sigma-UB (two-track geometry diagnostics). At each evaluated checkpoint (generation $g$ in S1/S2), we generate one response per frozen prompt (fixed decoding), embed responses with a frozen sentence encoder, and compute the two SIGMA-UB diagnostics from Section 3.4 (Algorithm 1) on a fixed observed index set (shared across checkpoints within a run). We report baseline drifts relative to the base checkpoint ($g=0$): (i) Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$ and (ii) Track I $\Delta G_{\mathrm{KF}}(\delta)$.
Interpretation. More negative drift indicates stronger contraction of the observed sub-Gram spectrum. Track II is the sensitive trend probe (observed-spectrum contraction); Track I is a conservative guardrail and may be loose when dominated by the worst-case tail-energy budget. We therefore interpret the joint behavior of both tracks as the primary diagnostic signal.
Surface-form proxies (non-definitional). To contextualize geometry drift, we also compute lightweight surface-form diversity proxies (distinct-2 and hashed $n$-gram HHI) on the same model outputs used for SIGMA-UB. Full proxy reporting is deferred to Appendix C.4.
4.5 Main-Text Reporting: Two Plots and One Table (FIN)
To keep the main text minimal and directly comparable across settings, we report: (i) two plots—one for S1 and one for S2—each showing both SIGMA-UB tracks, and (ii) one compact table summarizing S1 vs. S2 via final drift and an OLS slope over generations. All underlying per-generation values used to render the plots and compute the summary scalars are provided in Appendix C.3.
@@ 4.5 Main-Text Reporting: Two Plots and One Table (FIN)
Figure 1: FIN, Setting S1 (restart-from-base): drift vs. generation under both SIGMA-UB tracks. Top: Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$. Bottom: Track I $\Delta G_{\mathrm{KF}}(\delta)$. Values are baseline drifts relative to the base checkpoint ($g=0$). S1 isolates data recursion only: synthetic text is regenerated each generation, but training restarts from the same base checkpoint. Appendix C.3 reports the exact per-generation table used to render this figure.
@@
Figure 2: FIN, Setting S2 (true recursion): drift vs. generation under both SIGMA-UB tracks. Top: Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$. Bottom: Track I $\Delta G_{\mathrm{KF}}(\delta)$. Values are baseline drifts relative to the base checkpoint ($g=0$). S2 compounds data recursion + weight recursion: each generation regenerates synthetic text and continues training from the previous generation weights. Appendix C.3 reports the exact per-generation table used to render this figure.
Setting $g_{\max}$ $\Delta U_{\mathrm{LLN,cov}}$ (final) slopeLLN,cov $\Delta G_{\mathrm{KF}}$ (final) slopeKF Track agreement?
@@
S2 50 -1536.562 -42.621 -2.426 2.350 no (Track II collapses; Track I does not)
Table 1: FIN summary statistics (S1 vs. S2). We summarize each drift curve using: (i) final drift at the last available generation $g_{\max}$, and (ii) an ordinary least-squares slope of drift versus generation (more negative implies faster degradation over the observed window; slopes are fit on the full drift series including the g=0g{=}0 baseline point). Track agreement reports whether both tracks trend negative (qualitative agreement) or whether Track II indicates strong contraction without a matching Track I response. Appendix C.3 provides the per-generation tables used to compute these scalars.
Result (S1 vs. S2). On FIN, restart-from-base recursion (S1) shows a mild but consistently negative drift under both tracks by $g_{\max}=50$ (final drifts $\approx-151$ and $\approx-142$), consistent with gradual contraction driven by the evolving synthetic pool. Under true recursion (S2), Track II exhibits a much larger negative drift (final $\approx-1537$; slope $\approx-42.6$), indicating substantially accelerated contraction when weights are carried across generations. In contrast, Track I remains near-zero and can even trend slightly positive over the same window, reflecting that the KF envelope is governed by a worst-case tail-energy budget rather than the observed spectrum itself; this Track II/Track I separation is an intended diagnostic feature of Sigma-UB.

### v2

# Sigma: Scalable Spectral Insights for LLM Model Collapse

@@ 1 Introduction
LLM Model Collapse (Shumailov et al., 2024a, b) refers to the phenomenon where Large Language Models (LLMs) progressively degrade towards a degenerate state. This state arises as a result of recursive training, where models are trained on iterations of data generated by other models, often previous versions of themselves. We term this model-generated output Synthetic Data, in contrast to organic data produced by human users.
@@
In the current landscape of AI research, the demand for reliable and high-quality training data has never been higher. The scale of the latest generation of LLMs is increasing at such a staggering speed that the natural generation of internet content by humans cannot keep pace. Consequently, utilizing synthetic data has become a vital, if not inevitable, step in LLM training. While this is a logical progression, it carries a significant caveat. Without the infusion of new data produced by human users, model collapse occurs as the model iterates on its own probability distribution. Ultimately, the model may converge to a state characterized by low variance and degraded representational quality (Shumailov et al., 2024a, b).
@@
Therefore, it is imperative to conduct a rigorous mathematical examination of LLM model collapse. What are the underlying mechanisms at the mathematical level? How can this degradation be tracked and quantified throughout the training process? Furthermore, how can we leverage these metrics to gain deeper insights into LLM dynamics and guide the development of more efficient training architectures? This paper aims to answer these questions.
@@
We introduce Sigma (Spectral Inequalities for Gram Matrix Analysis), a novel framework that quantifies and benchmarks model collapse by utilizing the spectral bounds of the Gram matrix of LLM embedding vectors. The Gram matrix is a classical mathematical concept named after Jørgen Pedersen Gram. It has a wide range of applications in fields such as Riemannian geometry and control theory, and has found critical utility in modern applied mathematics and computer science domains, including the finite element method (Lanckriet et al., 2004). It is also widely used in classical NLP as an intuitive method to encode variable quantities of natural language embeddings into a fixed-dimensional matrix representation(Lodhi et al., 2002; Levy and Goldberg, 2014). As we enter the era of LLMs, the Gram matrix remains a vital tool, as it captures the intrinsic geometry and distributional properties of the data we seek to analyze. Sigma revisits this fundamental operator to quantify the modern phenomenon of representation collapse.
@@
Inevitably, the spectrum of the Gram matrix becomes the key to understanding LLM model collapse. We develop our theory and framework around this central mathematical object, allowing us to monitor, estimate, and predict the behavior of LLMs regarding model collapse through training iterations. Crucially, our method utilizes both deterministic and stochastic bounds to function effectively even with partially observed data, ensuring our framework remains scalable for large-scale models.
@@ 2.1 Related Work
The risk of training generative models on their own outputs has been formalized as model collapse, where repeated self-training progressively erodes distributional support, particularly in the tails, and can culminate in highly generic or degenerate generations (Shumailov et al., 2024a, b).
Related phenomena have been studied under different labels and experimental protocols, including Model Autophagy Disorder (MAD), which emphasizes precision–recall tradeoffs and the role of “fresh” real data injections across generations (Alemohammad et al., 2024). Recent theory sharpens these concerns by distinguishing regimes where degradation is transient versus strong and asymptotic, even when synthetic data is mixed with a constant fraction of real data (Dohmatob et al., 2025).
Complementary analyses interpret recursion as an entropy- or variance-contracting process and propose tail-centric quantities such as time-to-forget and covariance shrinkage to explain why rare events vanish first (Seddik et al., 2024). Other lines formalize stability of self-consuming loops via recursion-stability notions, connecting collapse to unbounded generalization degradation across generations (Yoon et al., 2025).
Finally, practical collapse signals in text include over-concentration of surface features (e.g., $n$-grams), emphasizing the need for diagnostics that are sensitive to diversity loss beyond average quality metrics (Zhu et al., 2025). Most prior work studies recursion within a single-model loop (or a fixed teacher–student chain), whereas real-world synthetic data arises from multi-model, retrieval-mediated ecosystems. Building on this black-box perspective, recent work develops statistically grounded uncertainty and calibration techniques tailored to LLM-as-a-Judge systems (Pang et al., 2025b, a), and studies related dynamics that can shift judge behavior over time (Wang et al., 2025), where LLM Web Dynamics (LWD) explicitly targeting this setting by modeling a network of LLMs coupled through a retrieval-augmented substrate, and analyzes convergence patterns under such web-like feedback.
@@ 3.3 Error Estimation
Naturally, we will be interested in how much error does the theorems in the previous section introduce. In fact, theorem 2 provides extra insights for the tightness of inequalities in 1. Let $C$ be the covariance matrix defined in theorem 2. We know that for large $n_{A},n_{k}$, $\det(G^{(k)})\approx\left(n_{k}/n_{A}\right)^{m}\det(G_{A}^{(k)})\approx n^{m}\det(C)$, we can define the asymptotic Overestimation Ratio $\mathcal{R}$ as the ratio of the Weyl bound to the stochastic expected value:
@@ 3.4 Sigma-UB Collapse Monitoring
Section 3.2 provides two complementary routes to control the full Gram determinant: a distribution-free deterministic bound (Theorem 1) and a stationary/i.i.d. scaling law (Theorem 2). However, the raw statements are not directly usable and practical because (i) the deterministic bound depends on the unobserved quantity $\beta_{k}=\lambda_{\max}(G_{B}^{(k)})$, and (ii) the stochastic scaling contains unknown constants and is meaningful primarily up to trends.
Our goal in Sigma-UB is therefore operational: to design online diagnostics that are fully computable from the observed summary statistics $\big(G_{A}^{(k)},\,n_{A},\,n_{k}\big)$, while remaining analytically trackable across checkpoints. This bridge is a central strength of the framework: even with partial Gram access (only the sub-Gram $G_{A}^{(k)}$ plus sample counts), we can still make principled inferences about the full Gram geometry, with uncertainty explicitly controlled via our deterministic envelope and stochastic scaling with error guarantees.
3.4.1 Metrics: two theorem-to-metric bridges
The geometric quantity we ultimately care about is the log determinant induced by the full Gram matrix,
$L^{(k)}(\delta)\;:=\;\log\det\!\big(G^{(k)}+\delta I_{m}\big),$
where we fix a small $\delta>0$ and reuse the same $\delta$ across generations. The $\delta I_{m}$ shift makes the log-determinant well-defined under rank deficiency and stabilizes numerics near collapse (Appendix B). The difficulty lies in the fact that we never explicitly construct $G^{(k)}$; instead, we only compute the observed sub-Gram matrix $G_{A}^{(k)}=A^{(k)}(A^{(k)})^{\top}$. Our framework, however, provides a principled way to interpret and infer properties of the full Gram matrix.
Bridge I (Theorem 1): a computable deterministic envelope.
Theorem 1 implies that the unobserved component can only inflate the spectrum, yielding a one-sided control of the form
$G^{(k)}\;\preceq\;G_{A}^{(k)}+\beta_{k}I_{m},$ (3.15)
which implies
$L^{(k)}(\delta)\;\leq\;\log\det\!\big(G_{A}^{(k)}+(\beta_{k}+\delta)I_{m}\big).$ (3.16)
The obstacle is that $\beta_{k}=\lambda_{\max}(G_{B}^{(k)})$ is unobservable.
To remove $\beta_{k}$, we enforce a standard preprocessing/normalization so that every embedding column satisfies $\|v_{j}^{(k)}\|_{2}^{2}\leq\rho$. With $n_{2}:=n_{k}-n_{A}$ unseen columns, a worst-case tail-energy budget gives the computable bound (Appendix B)
@@ 3.4.1 Metrics: two theorem-to-metric bridges
Substituting $\widehat{\beta}_{k}$ yields a valid one-sided envelope for $L^{(k)}(\delta)$. Finally, because the envelope level is dominated by the isotropic shift $(\widehat{\beta}_{k}+\delta)I_{m}$, we normalize it to isolate the gain contributed by the observed spectrum:
@@
$\qquad\quad\qquad-m\log(\widehat{\beta}_{k}+\delta).$ (3.18)
Equivalently,
$\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)=$ $\log\det\!\left(I_{m}+\frac{1}{\widehat{\beta}_{k}+\delta}G_{A}^{(k)}\right)$
$=$ $\sum_{i=1}^{m}\log\!\left(1+\frac{\lambda_{i}(G_{A}^{(k)})}{\widehat{\beta}_{k}+\delta}\right),$ (3.19)
so $\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ decreases when the observed eigenvalues contract. This metric is deterministic and conservative: it requires no distributional assumptions, and its conservativeness is governed solely by an explicit upper bound on the unobserved tail spectral mass.
Bridge II (Theorem 2): size correction and covariance normalization.
Under the stationary/i.i.d. assumption of Theorem 2, the full Gram determinant scales approximately like
$\det(G^{(k)})\approx\left(\frac{n_{k}}{n_{A}}\right)^{m}\det(G_{A}^{(k)}),$
i.e., the log-determinant shifts by the purely geometric term $m\log(n_{k}/n_{A})$. Moreover, when $G_{A}^{(k)}\approx n_{A}C$ for a stable covariance $C$, we have $\log\det(G_{A}^{(k)})\approx m\log n_{A}+\log\det(C)$. This motivates a size-invariant proxy for the underlying covariance geometry:
@@
Intuitively, $\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)$ tracks $\log\det(C)$: it removes the trivial $m\log n_{A}$ growth, so sustained negative drift reflects a genuine contraction of the representation geometry rather than a change in sample size. While the absolute scaling in Theorem 2 involves an unknown constant, this constant cancels when we monitor trends.
To remove checkpoint-independent constants and emphasize temporal contraction, we report drift relative to a baseline checkpoint $k=0$:
@@
3.4.2 Online monitoring procedure
To monitor the health of the representation space continuously during training, we utilize a partial observation strategy. At each checkpoint $k$, we compute the sub-Gram matrix $G_{A}^{(k)}$ on a fixed observed block size $n_{A}>m$. This allows us to evaluate both the Track-I (deterministic) and Track-II (stochastic) diagnostics using only the summary statistics $\big(G_{A}^{(k)},n_{A},n_{k}\big)$, avoiding the prohibitive cost of full eigendecomposition.
Algorithm 1 details the computational steps. In practice, the $\log\det(\cdot)$ terms are computed efficiently via Cholesky decomposition on the shifted positive semi-definite (PSD) matrices (see Appendix B).
@@ 3.4.2 Online monitoring procedure
Both metrics are designed to decrease as the eigenvalues of $G_{A}^{(k)}$ contract, becoming increasingly sensitive as the spectrum compresses toward the regularization floor $\delta$. A sustained negative drift in either $\Delta\mathcal{G}_{\mathrm{KF}}^{(k)}$ or $\Delta\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}$ signals the onset of collapse, while saturation at the floor suggests the model has reached a late-stage, rank-deficient state.
The dual-track design provides complementary theoretical insights. $\Delta\mathcal{G}_{\mathrm{KF}}$ acts as a conservative envelope. Being deterministic, it guarantees a valid upper bound on collapse given a fixed tail-energy budget. However, its theoretical sharpness degrades when the embedding space is highly anisotropic—i.e., dominated by a few prominent eigen-directions. In such cases, the scalar energy cap $\widehat{\beta}_{k}$ significantly overestimates the unobserved tail contribution, potentially masking early signs of contraction.
Conversely, $\Delta\mathcal{U}_{\mathrm{LLN,cov}}$ serves as a sensitive probe. By leveraging the Law of Large Numbers (LLN), it corrects for dataset size to estimate the intrinsic geometric density. Its limitation is statistical: if the underlying data distribution is highly complex or heavy-tailed relative to the block size $n_{A}$, the LLN convergence may be slow, introducing estimation variance. Consequently, a divergence where Track-II plummets while Track-I remains stable is often an informative signature of early collapse—indicating that the geometry is contracting (visible to the sensitive probe) but has not yet violated the worst-case energy budget (Track-I).
@@ 4 Experiments
We empirically stress-test recursive synthetic self-training under stable evaluation conditions to pinpoint where collapse may arise in the data cycling loop. The central question is not whether recursion can degrade outputs, but whether the observed geometric contraction sufficiently represents and tracks the collapse mechanism across protocols. Throughout, we track collapse using the two Sigma-UB diagnostics introduced in Section 3.4, computed on a frozen evaluation protocol so that measured drift is attributable to model evolution rather than evaluation drift.
All experiments in the main text are run and trained on a single controlled benchmark dataset (TECH) from DecayBench. We focus on TECH (software & technical domain contextual data) for two practical reasons: (i) it provides a large, heterogeneous set of prose spanning conceptual guides, tutorials, and operational documentation; and (ii)it provides a fixed evaluation prompt bank that can be reused unchanged across checkpoints, ensuring the distribution does not drift due to uncontrolled prompt-relevent variables. We construct a real TECH corpus $R$ as a collection of $N=1000$ fixed-length 64-token contexts. Evaluation uses a frozen TECH prompt bank $P$ with stable prompt IDs and a fixed bucket taxonomy (Creative, Divergent, Analogy/ELI5, What-if, Neutral). Every checkpoint is evaluated on the same prompt IDs under fixed decoding and embedded with a frozen sentence encoder. This prevents “evaluation drift” (where changing prompts/decoding would create apparent metric changes unrelated to the model). To ensure we can form stable Gram matrices with embedding dimension $m=384$, we generate multiple stochastic responses per prompt at each checkpoint; the full evaluation protocol and seed schedule are specified in Appendix C.3.
@@
At generation $g\geq 1$, we build a fresh synthetic pool $S^{(g)}$ by prompting the previous-generation model with each real context $x\in R$. This produces one synthetic continuation per training context, so $|S^{(g)}|=|R|$ and token budgets remain comparable across generations. The training data at generation $g$ is a constant-budget mixture of both synthetic and real date: a fraction $(1-\alpha)$ is drawn from $S^{(g)}$ and a fraction $\alpha$ from $R$, with the total number of 64-token training blocks held fixed to $|R|$.
Two recursion settings. We study two matched settings that share similar regeneration protocol and mixture rule and differ only in whether model weights are reset between generations.
@@
S1: restart-from-base (data recursion only). Each generation regenerates a fresh synthetic data pool, but training always restarts from the same base model at initial state:
$M^{(g)}\leftarrow\mathrm{Finetune}\big(M^{(0)},\,D^{(g)}(\alpha)\big).$ (4.1)
S1 isolates the impact of increasing synthetic data share in LLM training: while the synthetic data distribution shifts across generations, there is no cumulative weight drift.
@@
S2: true recursion (data + weight recursion). In this setting, we do not restart from a fixed base checkpoint. Instead, we carry the model weights forward across generations and finetune iteratively under the same regeneration and mixture rule:
$M^{(g)}\leftarrow\mathrm{Finetune}\big(M^{(g-1)},\,D^{(g)}(\alpha)\big).$ (4.2)
S2 is included because it matches the real-world self-consuming pipeline: each generation both produces the next synthetic data and inherits the parameters shaped by earlier generations. This induces a feedback loop that can magnify small changes over time. In addition, S2 and S1 comparision isolates the incremental impact of weight recursion: it tests whether repeatedly updating parameters on recursively generated data produces extra (and possibly accelerating) contraction, beyond the drift caused by data feedback alone under restart-from-base training.
@@
Figure 1: TECH, Setting S1 (restart-from-base): drift vs. generation under both SIGMA-UB tracks. Top: Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$. Bottom: Track I $\Delta G_{\mathrm{KF}}(\delta)$. Values are baseline drifts relative to the base checkpoint ($g=0$). S1 isolates data recursion only: synthetic text is regenerated each generation, but training restarts from the same base checkpoint. Appendix C.4 reports the exact per-generation table used to render this figure.
@@
Figure 2: TECH, Setting S2 (true recursion): drift vs. generation under both SIGMA-UB tracks. Top: Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$. Bottom: Track I $\Delta G_{\mathrm{KF}}(\delta)$. Values are baseline drifts relative to the base checkpoint ($g=0$). S2 compounds data recursion + weight recursion: each generation regenerates synthetic text and continues training from the previous generation weights. Appendix C.4 reports the exact per-generation table used to render this figure.
Table 1: TECH summary statistics (S1 vs S2). Final drifts and OLS slopes are computed from the plotted drift curves.
Setting g_max DeltaU_final slope_LLN,cov DeltaG_final slope_KF Track agreement?
@@
S2 50 -1536.562 -42.621 -2.426 2.350 no (Track~II collapses; Track~I does not)
Result (TECH; S1 vs. S2). Restart-from-base recursion (S1) exhibits a mild but consistently negative drift under both tracks through $g_{\max}=50$ (final drifts $\approx-151$ and $\approx-142$), consistent with gradual contraction induced by the evolving synthetic pool even when weights are reset each generation. Under true recursion (S2), Track II contracts by more than an order of magnitude (final $\approx-1537$; slope $\approx-42.6$), indicating substantially accelerated observed-spectrum contraction when weights are carried across generations. Track I, in contrast, remains near zero and can even trend slightly positive over the same window; this reflects that the KF envelope is governed by a worst-case tail-energy budget rather than the observed sub-spectrum itself. Accordingly, Track II/Track I separation is an intended diagnostic signature of Sigma-UB and is operationally useful for flagging “early collapse” regimes where the observed spectrum has contracted sharply while the conservative envelope remains permissive.
5 Conclusion
In this work, we established Sigma as a scalable framework for monitoring the onset of model collapse without the prohibitive cost of full eigen-decomposition. By deriving rigorous spectral inequalities, we demonstrated that the health of an LLM’s representation space can be accurately estimated using only a sub-sample of the embedding Gram matrix, effectively decoupling monitoring costs from the total corpus size.
Our empirical analysis of recursive loops reveals a critical distinction between data-only recursion (S1) and true recursion with weight carryover (S2). We showed that while data recursion induces gradual drift, the compounding effect of weight updates accelerates geometric contraction by orders of magnitude. Crucially, the divergence between our conservative deterministic envelope (Track I) and the sensitive stochastic scaling metric (Track II) serves as a diagnostic signature for early-stage collapse, often preceding surface-level repetition artifacts.
Future Work
This paper marks a rigorous first step in quantifying the static geometry of model collapse. However, the recursive training loop is inherently a dynamical process—a trajectory of distributions evolving over time. Our future work aims to transition from monitoring states to predicting dynamics by viewing the LLM training loop through the lens of Koopman Operator Theory (Brunton et al., 2022).
While the state-space evolution of model parameters is highly non-linear and opaque, the Koopman operator lifts these dynamics into an infinite-dimensional function space where the evolution becomes linear. This powerful shift allows us to deploy the robust tools of functional analysis and ergodic theory to analyze the system’s spectral properties. By characterizing the eigenvalues of the Koopman operator associated with the recursive training map, we hope to rigorously predict asymptotic behaviors—such as the rate of convergence to degenerate fixed points—and identify stable “eigen-distributions” that resist collapse. This approach promises not only to demystify the black-box nature of iterative learning but also to bridge the gap between modern dynamical systems theory and the empirical realities of foundation model maintenance.